a carpenter has a wooden dowel that is 72 cm long. she wants to cut it into 2 pieces so that one is 5 times as long as the other. what are the lentghs of the two pieces?
Answer by rapaljer(4671)

Let x = smaller piece
5x= larger piece

x+5x=72
6x=72
x=12 cm smaller piece
5x=5(12)=60 cm larger piece

Check: 12 +60=72
It checks!

If you need additional help with Word Problems, check out my website by doing a "Bing" search for my last name "Rapalje". Look for "Rapalje Homepage" at the top of the search list, and choose the link on my homepage "Basic, Intermediate and College Algebra: One Step at a Time." Click on "Basic Algebra", and you will find a couple of sections near the end of Chapter 1 on "Word Problems". My non-traditional explanation should be easier to understand than your textbook!
R^2
Dr. Robert J. Rapalje, Retired
Seminole Community College
Altamonte Springs Campus
Florida